What is the largest integer n which makes n3 + 100 is divisible by n + 10


1
Expert's answer
2021-09-02T00:37:22-0400
"n^3+100=n^3+1000-900"

"=(n+10)(n^2-n+100)-900"

The term "(n+10)(n^2-n+100)" is divisible by "n+10."

The Divisors of 900 are as follows:

1, 2, 3, 4, 5, 6, 9, 10, 12, 15, 18, 20, 25, 30, 36, 45, 50, 60, 75, 90, 100, 150, 180, 225, 300, 450, and 900.

We take "n+10=900." Then "n=890."

The largest "n" which makes "n^3+100" is divisible by "n+10" is therefore "890."